please help me solve this problem

A rectangular garden is 15 ft longer than it is wide. Its area is 4500 feet squared. What are its dimensions?

Answer by Theo(13342)

L = length
W = width

L = W + 15

A = area = L*W

L*W = 4500

Since L = W + 15, this becomes:

(W + 15) * W = 4500

This becomes W^2 + 15*W = 4500

Subtract 4500 from both sides to get W^2 + 15*W - 4500 = 0

Factor to get (W - 60) * (W + 75) = 0

This makes W = 60 or W = -75

W can't be negative so W = 60 or not at all.

Take W = 60

This makes L = W + 15 = 75

60 * 75 = 4500 so the area is good.

L = 15 more than W so L and W are ok.

Your answer is that the dimensions are:

Length = 75 feet
Width = 60 feet
Area = 4500 square feet